# The emergence of inherently 9-dimensional one-loop couplings from T-duality

### v1

@@ 4 Conclusion
It is known in the literature that the S-duality of the 10-dimensional type IIB effective action at order $\alpha^{\prime 3}$ involves contributions from the perturbative sphere- and torus-levels, as well as non-perturbative effects, which combine to achieve $SL(2,\mathbb{Z})$ invariance [12]. The circular reduction of these combinations also results in invariant terms in nine dimensions.
In this paper, we have shown that purely stringy couplings, which arise only at loop-level due to winding momentum along the circle, produce inherently nine-dimensional couplings. These are not descended from the 10-dimensional effective action and are invariant under $SL(2,\mathbb{R})$ transformations on their own. In particular, while S-duality forbids a 10-dimensional Chern-Simons term in type IIB, the inherently nine-dimensional Chern-Simons term—produced by imposing T-duality on the KK reduction of the Chern-Simons term from 10-dimensional type IIA theory—is invariant under S-duality.
Moreover, we have observed that the inherently nine-dimensional effective action of type IIB in the gravity–radius–vector sector, produced by imposing T-duality on the KK reduction of the pure gravity couplings of type IIA theory, is also S-duality invariant.
@@
Our demonstration of S-duality invariance for the small-radius, nine-dimensional, one-loop couplings in type IIB theory relied on the assumption of a constant dilaton and vanishing RR fields. Leveraging this established symmetry, the RR and non-constant dilaton terms can be incorporated into the nine-dimensional type IIB framework. This, in turn, allows one to use T-duality to determine the corresponding one-loop couplings in nine-dimensional type IIA theory at the large-radius limit. These results can then be lifted to ten dimensions to find the type IIA one-loop couplings at order $\alpha^{\prime 3}$. Since the circular reduction of M-theory couplings at the eight-derivative order is expected to yield these specific type IIA couplings, it would be a compelling future direction to use this prescription to determine the precise form of the M-theory couplings at the eight-derivative order.
The construction of a covariant action for type IIB supergravity is obstructed in ten dimensions by the self-duality of the RR four-form, a constraint that can only be incorporated through auxiliary fields [48, 49, 50]. No such obstruction exists for type IIA supergravity. Given that T-duality acting on the KK reduction of type IIA couplings yields covariant type IIB couplings in nine dimensions, it is a promising research direction to use this prescription to derive the full covariant effective action for RR fields in nine-dimensional type IIB theory—an action that may not be uplifted to a covariant ten-dimensional one.

### v2

@@ 4 Conclusion
It is known in the literature that the S-duality of the 10-dimensional type IIB effective action at order $\alpha^{\prime 3}$ involves contributions from the perturbative sphere- and torus-levels, as well as non-perturbative effects, which combine to achieve $SL(2,\mathbb{Z})$ invariance [12]. The circular reduction of these combinations also results in invariant terms in nine dimensions [13].
In this paper, we have shown that purely stringy couplings, which arise at loop level from winding modes on the circle, correspond to inherently nine-dimensional terms in the effective action. These couplings are not descended from ten-dimensional terms and are separately invariant under the continuous $SL(2,\mathbb{R})$ symmetry. In particular, we have derived two distinct classes of such terms: the inherently nine-dimensional Chern–Simons coupling, obtained by applying T-duality to the KK reduction of the type IIA Chern–Simons term; and the inherently nine-dimensional couplings in the metric–radion–vector sector of type IIB theory, obtained by imposing T-duality on the KK reduction of the pure gravity couplings of type IIA theory. Both sets of couplings are invariant under $SL(2,\mathbb{R})$ transformations when the dilaton is held constant and the RR fields are set to zero.
@@
We demonstrate S-duality for the one-loop couplings in nine-dimensional type IIB theory at small radius, under the assumption of a constant dilaton and vanishing RR fields. The first step is to extend this result by incorporating RR fields and a non-constant dilaton into the nine-dimensional type IIB framework, maintaining S-duality invariance. We then apply T-duality to this generalized action to derive the corresponding one-loop couplings in nine-dimensional type IIA theory at large radius. Lifting these results to ten dimensions yields the one-loop $\alpha^{\prime 3}$ terms in type IIA theory. These terms should originate from the circle reduction of eleven-dimensional M-theory couplings, which relates them directly to M-theory. In line with this expectation, recent results in [35, 46] have identified related couplings involving RR fields and the dilaton that match the M-theory Chern-Simons term and pure gravity couplings. A compelling future check would be to explicitly show that the circular reduction of the type IIA couplings found in these references indeed exhibits the expected $SL(2,\mathbb{R})$ symmetry in nine dimensions after a T-duality transformation.
Using the fact that dimensional reduction of M-theory covariant couplings on $S^{1}\times S^{1}$ produces Lorentz-invariant couplings in nine-dimensional type IIA theory with a global $O(2)$ symmetry [50], one expects that any eleventh-dimensional couplings at the eighth-derivative order would yield corresponding couplings in type IIA theory possessing the same $O(2)$ symmetry. After applying T-duality, these would produce couplings in type IIB theory invariant under $SL(2,\mathbb{R})$ transformations. Consequently, this nine-dimensional symmetry alone cannot serve as a constraint to uniquely determine the M-theory couplings.